Rylee88:

Which expression is equivalent to 3(8 + 7)?

AZ:

You need to know the distributive property \( a(b+c) = ab + ac\) 

AZ:

You multiply that number that is outside the parenthesis with each term inside the parenthesis

Rylee88:

So I do x3?

AZ:

yes, with the 8 and the 7 and then you add them do you want to give us the answer choices?

Rylee88:

Yes there, 24 + 7 24 + 21 11 + 10 11 + 7

AZ:

Okay so what do you get when you multiply 8 and 7 with 3? 3(8 + 7) = (3*8) + (3*7)

Rylee88:

3*8=23? 3*7=20?

AZ:

how??? just use a calculator

Rylee88:

lol ok

Rylee88:

24 and 21

AZ:

there you go

Rylee88:

Ok thank You :)

AZ:

You're welcome